OpenStudy (anonymous):

The rate of change of y with respect to t is 3 times the value of the quantity 2 less than y. Find an equation for y, given that y = 212 when t = 0. You get: A. y=212e^3t-2 B. y=214e^3t-2 C. y=210e^3t+2 D. y=212e^3t+2 E. y=210e^3t-2

OpenStudy (anonymous):

\[ \frac{dy}{dt} = 3 \times (y-2) \]\[\int \frac{dy}{y-2} = 3 \int dt\]\[\ln|y-2| = 3t + c\]\[y - 2 = e^{3t+c}\]\[y-2=c_1e^{3t}\] Using the initial conditions we have: \[212 - 2 = c_1e^{(0)}\]\[c_1 = 210\] Substituting back in the equation, we have: \[y = 210e^{3t} + 2\] Option C